### v1

# AI Rules? Characterizing Reddit Community Policies Towards AI-Generated Content

Abstract.
How are Reddit communities responding to AI-generated content? We explored this question through a large-scale analysis of subreddit community rules and their change over time. We collected the metadata and community rules for over $300,000$ public subreddits and measured the prevalence of rules governing AI. We labeled subreddits and AI rules according to existing taxonomies from the HCI literature and a new taxonomy we developed specific to AI rules. While rules about AI are still relatively uncommon, the number of subreddits with these rules almost doubled over the course of a year. AI rules are also more common in larger subreddits and communities focused on art or celebrity topics, and less common in those focused on social support. These rules often focus on AI images and evoke, as justification, concerns about quality and authenticity. Overall, our findings illustrate the emergence of varied concerns about AI, in different community contexts.
Online Communities, Reddit, Rules, AIGC, Generative AI, Governance, AI-Generated Content, Moderation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYccs: Human-centered computing Empirical studies in collaborative and social computing
1. Introduction
The emergence of mainstream generative AI tools capable of producing compelling text, images, and videos—collectively referred to as AI-Generated Content (AIGC)—poses a unique challenge for online platforms that are dependent on User-Generated Content. Many such platforms have enacted top-down changes to respond to this new technology, with some deciding to ban it (Makyen, 2023), while others are encouraging its use or even incorporating generative AI features into their user interfaces (Lunden, 2023; Mehta, 2023). Other platforms have declined to make top-down changes and have instead left it up to their users to come to their own positions on this technology. One such platform is Reddit, a social sharing and news aggregation site, where individual communities, known as subreddits, are free to enact their own rules and policies (Fiesler et al., 2018; Proferes et al., 2021; Singer et al., 2014; Gilbert, 2013). This decentralized moderation approach allows communities to craft policies towards generative AI that are contextually-relevant: the policies align with how communities see the technology being used, and whether or not that use aligns with their values. Decentralized moderation also provides a rich source of data through which HCI researchers interested in helping online communities navigate sudden technological change can better understand the range of attitudes and stances that exist across communities who vary in size, purpose, and values.
Findings from past HCI studies have emphasized the important role that explicit and consistent community rules play in helping online communities navigate moments of intense change (Kiene et al., 2016). These rules interact with other forms of governance (Lessig, [n. d.]), such as norms (Chandrasekharan et al., 2018) and technical affordances (Jhaver et al., 2019), to encourage certain behaviors and discourage others (Kiesler et al., 2012).
Of course, generative AI and AIGC may impact the rules (Makyen, 2023), dynamics (Hua et al., 2024; Fiesler, 2024), and governance (Cao et al., 2023) strategies of online communities in complex ways. While some recent research has explored how generative AI might help online communities deal with problematic content (Agarwal et al., 2024; Kumar et al., 2024), other work has documented the ability of AIGC to pollute these same communities with deceptive information (Park et al., 2024). Recent work by Lloyd et al. (Lloyd et al., 2023) used interviews with Reddit moderators to explore the particular challenges of governing AI use in online communities and found that communities are enacting explicit rules about AIGC in order to clarify norms. Lloyd et al.’s work focuses on the perspectives and experiences of their interviewees and raises the question of how the range of concerns about AIGC that the moderators voice, both practical and ideological, generalize across the wide variety of communities on Reddit.
This paper addresses that question through a large-scale quantitative analysis of community rules on Reddit that govern AI usage. Our approach allows us to paint a fuller picture for the HCI community by complementing Lloyd et al.’s qualitative findings with an analysis of platform-wide patterns and trends, while simultaneously providing deeper insight into the specific types of AI rules that communities are creating. We build on quantitative techniques developed by Fiesler et al. (2018) to study the general ecosystem of community rules on Reddit and adapt their methodology to specifically investigate the emergence of community rules that govern AI use. We use the same community topic and rule taxonomies as Fiesler et al. so that we can compare our analysis of AI rules to their analysis of Reddit rules in general. We address the following research questions:
  • RQ1: How common are subreddit rules governing the use of AI and how has this changed over the last year?
  • RQ2: What types of communities are more likely to have these rules?
  • RQ3: What types of rules exist for governing the use of AI?
To answer these questions we crawled over $300,000$ public subreddits twice, during the summers of 2023 and 2024, in order to build a dataset of $n=80,290$ English language subreddits with rules. Our inquiry into RQ1 finds that while only $3\%$ of subreddits with rules have rules about AI, the number of subreddits with AI rules has almost doubled between the time of our crawls. More specifically, larger subreddits are much more likely to have AI rules: for example, of the top $1\%$ largest subreddits with rules, almost $19\%$ had AI rules in our most recent crawl. To explore RQ2, we perform subreddit-level analysis on all subreddits in our sample with rules. We apply a “Topic” label to these subreddits using a taxonomy from Fiesler et al. (Fiesler et al., 2018) and “Community Archetype” labels using a taxonomy from Prinster et al. (Prinster et al., 2024). We use quantitative methods to explore the relationship between these labels and the presence of AI rules and find these rules to be most common in communities with Art and Celebrity topics, as well as those that fit the Content Generation Community Archetype. To explore RQ3, we perform rule-level analysis on all the rules in our sample that we identified as governing AI use. We apply an existing rule taxonomy from Fiesler et al. (Fiesler et al., 2018) and additionally produce our own set of emergent labels to describe meaningful dimensions of variation in both the requirements imposed by these rules and the language through which they frame AI usage. We find rules about AI images to be more common than rules about other types of AIGC and see quality and authenticity as the most frequently expressed concerns.
We compare these findings to the results of related work and discuss implications for platform designers. Our findings identify some common community concerns about quality and authenticity, especially related to images, that may be addressed through thoughtful design. Our findings also show, though, that communities of different types take varied approaches towards regulating AI use, and suggest that design solutions meant to help communities address these challenges must be sensitive to community context.
2. Related Work
Our study is motivated by two bodies of work from the HCI literature. The first set of related work includes empirical research into online communities’ rules, norms, and values. This work demonstrates that online communities hold a wide range of values and emphasizes that rules and norms are always context-dependent. The second body of work concerns current and potential future impacts of generative AI on online communities. This research shows that communities are already grappling with the changes brought by generative AI and are adapting their policies and practices in response.
2.1. Online Communities’ Rules, Norms, and Values
Online communities play an important role in social life (Preece, 1998), and occupy a complicated position as a site of nuanced interactions between people, technology, and policy (Jackson et al., 2014). This importance and complexity have made online communities a frequent topic of study in the HCI literature (Preece, 2000; Kraut and Resnick, 2012). HCI scholars have explored the variety of purposes that these communities can serve and how platform design choices may influence their success. For example, Preece introduced the idea of Empathic Communities (Preece, 1999), in which people come together to share and receive support, and suggested platform design guidelines to encourage sociality (Preece, 2001). In addition to encouraging pro-social behavior, a frequent topic of study has been how to discourage anti-social behavior (Kiesler et al., 2012; Lessig, [n. d.]). Much of this work has focused on the benefits (Grimmelmann, 2015) and challenges (Koshy et al., 2023) of effective and fair content moderation. The HCI community has been interested in top-down moderation (Gillespie, 2018), in which platforms set and enforce rules, but even more interested in bottom-up, or self-moderation, in which communities do (Seering, 2020). A major finding of this research is the importance of context sensitivity in moderation decisions, which self-moderation is often more attuned to. Taken together, this research suggests that studying self-moderation decisions may provide insight for design solutions that can promote community self-determination at a time of sudden technological change.
Reddit is a fruitful platform for a study of community self-moderation because it is home to online communities of various size and purpose. This heterogeneity, combined with its largely public and text-based nature, have made Reddit a frequent topic of study in online communities research (Proferes et al., 2021). Recent work has explored the values (Weld et al., 2024), norms (Chandrasekharan et al., 2018), and rules (Fiesler et al., 2018) of different communities across the platform, both via direct methods like surveys, and indirect methods such as analyzing the digital trace data produced in these communities. Reddit’s implementation of community self-moderation allows communities to write and make public their own community-enforced rules, which Fiesler et al. collect and characterize in a study that is the main methodological inspiration for our work (Fiesler et al., 2018). In their study, Fiesler et al. find subreddit size to be the main predictor of having rules, but note that even among subreddits of similar sizes rules are context-dependent, not uniform, and often uniquely tailored to the topic of the subreddit (e.g., art subreddits have more rules about copyright). The authors also highlight that setting norms and developing rules is fundamentally a social process. Overall, Fiesler et al. show the importance of community rules and the feasibility of using them as a data source to study attitudes and stances across a platform. Given the potential impact of AI on online communities, we use community rules as a starting point to study context-specific attitudes towards AIGC as they emerge across the platform.
2.2. Generative AI’s Impact on Online Communities
A growing body of research in HCI, Social Computing, and adjacent behavioral science fields has documented ways in which generative AI may impact online communities through increased deception, decreased trust, and new forms of harassment. Studies have demonstrated that AI can effectively deceive humans (Park et al., 2024) and have raised concerns that generated text could be weaponized for deception by either humans (Goldstein et al., 2023) or automated social media accounts (bots) (Menczer et al., 2023). Additionally, just the suspicion that other online community members are using AI may be enough to impact dynamics in online communities. Several studies in AI-Mediated Communication (AIMC) (Hancock et al., 2020) provide evidence that the perception that another online community member is using AI causes members to regard each other more negatively (Rae, 2024) and trust each other less (Jakesch et al., 2019). Finally, HCI research has long documented how online communities can be sites of harassment and abuse (Jiang et al., 2019; Park et al., 2022; Cheng et al., 2017; Han et al., 2023). The impact of AI-powered harassment on online communities, such as malicious synthetic, or deepfake, media (Farid, 2022) is much less studied (Umbach et al., 2024), but there is concern that this extremely harmful phenomenon may grow in prevalence alongside the popularity of generative AI tools (Rini and Cohen, 2022).
Many additional studies have started to document the impacts of AIGC on online communities “in the wild”. Recent work has identified undisclosed AI use in several settings, such as LLM-powered social bots on Twitter (Yang and Menczer, 2024), AIGC in fake online reviews (Oak and Shafiq, 2024), and spam content used for scams (DiResta and Goldstein, 2024). In addition to studies of deceptive use, other recent work has begun to explore how the disclosed use of AIGC may affect community dynamics, such as in art-sharing communities (Wei and Tyson, 2024). Generative AI can additionally impact online communities by drawing traffic away from communities and decreasing overall engagement: causal inference studies of the Q&A sites in the Stack Exchange network, where users gather to ask and answer questions about programming and related topics, have demonstrated that the launch of ChatGPT caused an overall decline in website visits and question volume (Burtch et al., 2024). The authors of this study contrast this decline with a measured null-effect on similar Reddit communities, which they suggest may avoid a decline in traffic because of their emphasis on socialization, rather than pure information exchange. Finally, recent work by Lloyd et al. used interviews to qualitatively explore the attitudes and experiences of Reddit moderators dealing with AIGC (Lloyd et al., 2023). This study interviewed moderators from subreddits of different sizes and topics and surfaces a variety of community objections to AIGC that range from practical to ideological. The work poses open questions about how prevalent these concerns are across the platform and suggests that larger trends could be explored quantitatively in future work. Our study fills that gap through a quantitative analysis that helps identify the most common concerns and the most impacted communities, so that the HCI community can begin the task of addressing these issues with context-specific solutions.
3. Methods
In order to perform a large-scale analysis of community rules governing the use of AI, we built four datasets consisting of the metadata of public, English-language subreddits through a process summarized in Figure 1. These datasets of subreddits and their associated rules include data for the same subreddits collected at two different points in time, one year apart. The datasets include: the Broad Subreddit Set ($n=264,592$) consisting of subreddits with and without rules, the Rules Subreddit Set ($n=80,290$) of only subreddits with explicit rules, the AI Rules Subreddit Set ($n=2,600$) of subreddits with rules governing the use of AI, and the AI Rules Set ($n=4,150$) of individual rules governing AI use. We then classified the AI Rules Set using emergent coding, through which we created a taxonomy of AI rules. We supplement the raw crawl data with additional information about subreddit and rule types through several rounds of labeling, using both human and LLM coders. We will make our datasets and labels public upon completion of the peer-review process.
Figure 1. The process for collecting subreddit metadata and building the datasets used in our analysis: the Broad Subreddit Set of English-language subreddits, the Rules Subreddit Set of subreddits with non-empty rules, the AI Rules Subreddit Set of subreddits with rules governing the use of AI, and the AI Rules Set of individual rules governing AI use.
[…]
A funnel diagram with five steps showing the filtering of subreddits based on various criteria: 1. Step 1: 337,399 subreddits, representing 100% of all subreddits seen during the 2023 crawl. 2. Step 2: 295,152 subreddits (87.5%), representing subreddits still accessible during the 2024 crawl. 3. Step 3: 264,592 subreddits (78.4%), filtered to only subreddits where English is the primary language. 4. Step 4: 80,290 subreddits (23.8%), filtered to subreddits with non-empty ”rules” during either the 2023 or 2024 crawl. 5. Step 5: 2,600 subreddits (0.771%), further filtered to subreddits with rules governing AI use during either crawl. The final output is 4,150 individual AI rules identified in these subreddits.
3.1. The Broad and Rules Subreddit Sets
To build the Broad Subreddit Set of $n=264,592$ public, English-language subreddits, we performed two automated crawls: one in the summer of 2023 and another one year later, in summer 2024. The first crawl collected metadata for all subreddits indexed by Reddit’s “Top Communities” list [^1] https://www.reddit.com/best/communities/1/, which includes public subreddits with more than ten subscribers. We used a web-scraping library [^2] https://docs.scrapy.org/en/latest/ to visit the webpage for each of these subreddits and extract relevant metadata from the page HTML. This first crawl scraped $n=337,399$ subreddits in total.
We performed our second crawl in the summer of 2024. We fetched up-to-date metadata for each subreddit seen during our first crawl. As our main goal here was understanding rule development over time, we removed from our sample $n=42,247$ subreddits that were inaccessible during this second crawl [^3] Subreddits could be inaccessible either because they no longer existed or were not public at the time that we attempted to crawl them.. We used an automated tool [^4] https://pypi.org/project/langdetect/ to determine the language of each subreddit based on the text in its description and rules metadata. We removed $n=30,560$ subreddits from our sample whose primary language was not English. This process left us with the $n=264,592$ of the Broad Subreddit Set which we used to understand platform-wide patterns in the data and factors that contribute to whether subreddits have rules.
Using this broad dataset, we constructed a second, smaller dataset: the Rules Subreddit Set ($n=80,290$). This dataset consists only of subreddits with non-empty rules metadata at the time of at least one of our crawls. We use this dataset to understand patterns specific to subreddits with rules. These two datasets are the focus of our descriptive analysis and were used to create two other datasets specific to AI rules, which are discussed in Sections 3.2.2 and 3.2.3.
3.2. Data Labeling
To answer our research questions about rules governing the use of AI, we needed labels that identified these rules and provided additional context about both the rules and the communities that used them. Through the process described below, we added Topic, Community Archetype and Has AI Rule? labels to each of the subreddits in the Rules Subreddit Set. From these labels we identified the subset of subreddits that had AI rules, which we refer to as the AI Rules Subreddit Set ($n=2,600$). The specific AI-related rules within these subreddits became our AI Rules Set ($n=4,150$) and were each labeled with Rule Type, AI Rule: Rationale, AI Rule: Stance, AI Rule: Medium, and AI Rule: Framing labels. Below, we describe the labels for subreddits and rules and the process through which we created them. The applied labels are summarized in Table 1 (subreddits) and Table 2 (rules).
Table 1. Labels applied to subreddits in the Rules Subreddit Set
Label Source Mutually Exclusive? Example Values
Topic Fiesler et al. Yes Art, Celebrity, Video Games
Community Archetype Prinster et al. No
Content Generation, Topical Q&A,
Social Support
Has AI Rule? New Yes Yes, No
Table 2. Labels applied to rules in the AI Rules Set
Label Source Mutually Exclusive? Example Values
Rule Type Fiesler et al. No Advertising & Commercialization, Prescriptive, Restrictive
AI Rule: Rationale New No Low Quality, Inaccurate, Inhuman
AI Rule: Stance New No Unqualified Ban, Qualified Ban, Disclosure Required
AI Rule: Medium New No Text, Images, Deepfakes
AI Rule: Framing New No Generate/Create, Edit/Assist/Enhance, Unspecified
3.2.1. Detecting Community Rules Governing AI Use
Community rules are often very long and complex, which made the Rules Subreddit Set too large to manually code for the presence of rules governing the use of AI. To get a rough sense of the prevalence of these rules, we first used a simple regex to search the rules in the dataset for AI-related keywords. This step found matches in the rules of less than $2\%$ of the subreddits. We then devised a protocol, inspired by recent HCI work using LLMs for qualitative labeling tasks (Choksi et al., 2024; Rao et al., 2024; Aubin Le Quéré et al., 2024), to detect AI-related rules across the entire dataset with more precision. The approach is based on comparing LLM-produced labels with those produced by human experts and iterating on the LLM prompt until desirable inter-rater reliability (IRR) is achieved.
[…]
We used these manually labeled rules as ground truth against which to evaluate the performance of an LLM coder, which we planned to use to label the rest of the dataset. We crafted a prompt that combined the human coder instructions with Chain-of-Thought prompting techniques (Wei et al., 2023), including the recommended eight few-shot, in-context examples, which we selected from the manually labeled sample. We produced LLM labels by using OpenAI’s API to prompt gpt-4o-mini, OpenAI’s current flagship small model, using this prompt and the same input data that were given to the human coders (a subreddit’s rules). We compared the LLM-produced labels with the human-produced ones and calculated an IRR of .7.7.7.7, which we deemed insufficient. To improve IRR we repeated this process of randomly sampling, manually coding, generating LLM labels, comparing the results, and updating the prompt using the examples that the model got wrong. We performed three iterations of this step, at which point the LLM labels had an IRR of .95.95.95.95 with the human labels. We then labeled the entire Rules Subreddit Set using OpenAI’s API and gpt-4o-mini. The final prompt used can be found in Appendix A.1. This process identified the AI Rules Subreddit Set ($n=2,600$) of subreddits with rules governing AI use.
3.2.2. Applying Subreddit Classification Labels
We next sought to label the subreddits in the Rules Subreddit Set with details about the type of community that each subreddit was home to. To this end, we devised a process to apply a Topic label using a taxonomy from Fiesler et al. (Fiesler et al., 2018) and Community Archetype labels using a taxonomy from Prinster et al. (Prinster et al., 2024). In applying our labels, we considered a subreddit’s name, topic, description, and rules metadata fields.
[…]
3.2.3. Identifying Emergent AI Rule Types
In order to better understand the types of rules that subreddits use to govern AI, we identified these rules in our sample and used an iterative inductive coding process to develop a descriptive taxonomy. To do this coding, we slightly modified the prompt developed and validated in Section 3.2.1 so that it would flag individual rules rather than a set of rules as being AI-related. Since the prompt required only a slight modification from the prior prompt, and since human coders would be manually reviewing the labels in the next step, we decided that we did not need to validate model performance before tagging the entire sample. We used this prompt to instruct gpt-4o-mini to label each of the $n=33,780$ individual rules in the AI Rules Subreddit Set, which produced the AI Rules Dataset ($n=4,150$) of individual rules about AI.
Next, two of the authors collaborated to identify emergent labels for the rules. The authors randomly selected $n=25$ rules from this dataset and independently performed affinity mapping (Harboe and Huang, 2015) by grouping rules that appeared similar into emergent categories. The authors then compared results and consolidated their categories into labels that best captured the elements of interest. Each of the two authors then independently applied this set of labels to another random sample of $n=25$ rules, once again grouping similar rules, including into new emergent groups where the existing labels did not suffice. The two authors repeated this process of randomly sampling, independently labeling, collaboratively comparing their results, and consolidating their labels, until no new labels emerged and they were confident that their taxonomy covered the relevant dimensions of variation in the dataset. The final label groupings were Rationale, Stance, Medium and Framing, which are listed alongside some of their possible labels in Table 2. We applied these labels to the entire AI Rules Dataset, as we describe next.
3.2.4. Applying Rule Classification Labels
We used LLMs to apply our emergent labels and an additional Rule Type label from Fiesler et al. (Fiesler et al., 2018) to the entire AI Rules Dataset. The two first authors collaboratively developed a codebook with instructions for applying these labels. Fiesler et al. does not provide codebook instructions, so the authors interpreted the examples presented in the study and came up with instructions for how to apply each of the $24$ possible labels 555“Rule Type” labels are not mutually exclusive, so we represented them as a series of binary labels, one for each possible Rule Type..
[…]
4. Findings
We group our findings by research question and present in sequence: (1) the prevalence of rules governing AI use, (2) the types of communities that have these rules, and (3) the types of rules that we find for governing AI use.
4.1. RQ1: How Common are Subreddit Rules Governing the Use of AI?
Table 3. Presence of Subreddit Community Rules Governing the Use of AI
N %
Rules Subreddit Set 80,290
w/ AI rules in 2023 crawl 1,411 1.76%
w/ AI rules in 2024 crawl 2,551 3.18%
YoY Increase 1,140 +81%
Top 1% Of Rules Subreddit Set by Subscriber Count 803
w/ AI rules in 2023 crawl 82 10.21%
w/ AI rules in 2024 crawl 152 18.93%
YoY Increase 70 +85%
Table 3 shows the results of our detection task described in Section 3.2.1. We find that during our first crawl in 2023, $n=1,411$ subreddits in the Rules Subreddit Set had rules governing the use of AI ($1.76\%$). By the time of our second crawl, this number increased by $81\%$ to $n=2,551$ ($3.18\%$). Rules governing the use of AI are more common in larger subreddits. For example, of the largest $1\%$ of subreddits in the Rules Subreddit Set by subscriber count, $10.21\%$ had AI rules during our first crawl. This percentage increased to $18.93\%$ in our second crawl. Figure 2 shows differences in the prevalence of AI rules for subreddits with different numbers of subscribers using data from the Rules Subreddit Sample. The figure contains two subplots which both have subscriber count split into deciles along the X axis (the labels represent the lower edge of the decile). The subplot on the left charts the percentage of subreddits with AI rules during our second crawl. The subplot on the right charts the percent increase in subreddits with AI rules between the two crawls. For example, the rightmost bar in the left subplot shows that in the largest decile of subreddits, about $11\%$ had rules about AI during the second crawl. The figure shows that subreddits with more subscribers are more likely to have AI rules, and that communities of all sizes saw increases in the prevalence of AI rules between our two crawls, though that increase was more prominent for larger communities.
Figure 2. Percentage of subreddits with AI rules, bucketed by decile based on subscriber count. The first chart shows the percentage of subreddits with AI rules at the time of our second crawl and the second chart shows the percent change between the two crawls.
[…]
The image contains two bar charts comparing the presence of AI-related rules in subreddits by subscriber count deciles:
Left Chart: ”Percentage of Subreddits With AI Rules”
This chart shows the percentage of subreddits with AI rules across different subscriber count deciles (measured by the lower edge of each decile). The subscriber counts (shown on the x-axis) are 11, 24, 48, 94, 188, 386, 850, 2155, 5978, and 22,345. The percentage of subreddits with AI rules steadily increases as subscriber count grows. The largest decile (22,345) has the highest percentage, above 10%, while the smallest decile (11) has a percentage near 0%. Right Chart: ”Increase in Subreddits With AI Rules”
This chart illustrates the year-over-year (YoY) percentage increase in the number of subreddits with AI rules, again grouped by subscriber count deciles. The subscriber counts (on the x-axis) are the same as the first chart: 11, 24, 48, 94, 188, 386, 850, 2155, 5978, and 22,345. The YoY increase is highest for the 94-subscriber decile, reaching almost 100%. Other significant increases are seen for the deciles of 188, 850, 5978, and 22,345, all of which have YoY increases around 80% to 90%.
4.2. RQ2: What Types of Communities Have Rules Governing AI Use?
In this section, we examine the characteristics of communities that have rules governing the use of AI. To do that, we compare the characteristics of two groups of subreddits in the Rules Subreddit Set: subreddits with rules about the use of AI, and those without such rules. We find that larger subreddits and those associated with Art or Celebrity topics are more likely to have explicit rules regulating AI.
Table 4. Differences between subreddits in the Rules Subreddit Set ($n=80,290$) with and without AI rules. For numerical subreddit characteristics, we show the average value across the subreddits in each group. For binary characteristics, we show the percentage of subreddits with this characteristic in each group and denote these percentages with (%). While we computed differences for all 28 topics, we show here only the ten topics with the largest absolute differences.
No AI Rule AI Rule Difference
Metadata Subscriber count 33488.25 274350.19 240861.94***
Subreddit Age (Years) 5.91 8.06 2.15***
Subreddit created after ChatGPT (%) 5.88 9.88 4.0***
Community Archetype Content generation (%) 51.97 70.56 18.59***
Affiliation with an entity (%) 56.23 61.58 5.36***
Topical question and answer (%) 7.24 6.74 -0.5
Learning and perspective broadening (%) 17.9 14.11 -3.79***
Social support (%) 8.39 2.43 -5.96***
Topic Celebrity (%) 2.95 20.42 17.48***
Art (%) 2.59 13.84 11.25***
Video Games (%) 5.78 7.64 1.86***
Politics (%) 2.32 0.51 -1.81***
Business and Organizations (%) 2.86 0.67 -2.2***
Games (%) 8.82 6.12 -2.7***
Local (%) 4.03 1.25 -2.77***
Hobby (%) 7.93 4.59 -3.34***
Support (%) 7.55 1.76 -5.78***
Humor (%) 10.27 3.45 -6.82***
Note: ${}^{*}p<0.1,^{**}p<0.05,^{***}p<0.01$
To identify subreddit characteristics that predict the existence of AI rules, we perform a difference in means test between subreddits with and without AI rules. Table 4 presents the results of the test, showing various subreddit characteristics that differ between these two subreddit types. For numerical subreddit characteristics like Subscriber Count and Age the table shows the average value across these two groups of subreddits. For binary characteristics, such as whether a subeddit was created after the launch of ChatGPT, or if it has a specific Community Archetype or Topic label, we show the percentage of subreddits in each group with this characteristic. To test for significance we use chi-squared tests for the binary characteristics and the non-parametric Mann-Whitney U test for numerical characteristics. We implement Bonferroni correction to adjust for multiple comparisons.
Table 4 exposes several key differences between subreddits with and without AI rules. Subreddits with AI rules tend to have more subscribers. These subreddits also tend to be newer and a higher percentage of subreddits with AI rules were created after ChatGPT was launched. We observe differences between the groups with respect to their association with the different Community Archetype labels. For example, there is a large and significant difference related to the Content Generation label; subreddits with this label are almost $20$ percentage points more likely to have AI rules. In contrast, subreddits of the Social Support Community Archetype are about $6$ percentage points less likely to have AI rules. Lastly, the Art and Celebrity topics are more common in subreddits with AI rules, encompassing about a third of all subreddits with AI rules compared to only $6\%$ of subreddits with no AI rules. These differences may indicate that communities focused on developing and sharing content are more likely to regulate AI usage in an attempt to maintain quality and standards, whereas communities that prioritize relationships among users are less concerned with the use of AI. We will further explore these implications in the discussion.
We additionally run several logistic regression (LR) models to validate our results. As a baseline, we repeat the analysis done in Fiesler et al. (2018) and use a LR model to predict the presence of any set of rules for subeddits in the Broad Subreddit Set, using as input variables the following metadata gathered in our second crawl: Subscriber Count [^6] We log transform Subscriber Count to account for the variable’s skewed distribution., Subreddit Age (in years), and a binary indicator of whether or not the subreddit was created after the launch of ChatGPT. We compare these results with a LR model that specifically predicts the presence of AI rules, using as input all of the subreddit characteristics shown in Table 4. Comparing the two regression results isolates the subreddit characteristics that are specifically predictive of AI rules, not rules in general. We find that the Art and Celebrity Topic labels are both predictive of AI rules ($\beta=1.49,p<0.01$ and $\beta=2.12,p<0.01$), supporting the analysis above. In terms of Community Archetype labels, the Content Generation label is predictive of AI rules ($\beta=0.71,p<0.01$), while the Social Support label is negatively associated with the presence of AI rules ($\beta=-0.45,p<0.01$). These regression results are in line with the findings from our difference in means test.
4.3. RQ3: What Types of Rules Exist for Governing AI?
While our previous analyses have helped explain what types of communities are likely to explicitly regulate AI, we also consider differences in how communities regulate AI. Drawing on the AI Rules Set ($n=4,150$), we produce frequencies of the different Rule Type labels described in Sections 3.2.3 and 3.2.4. Table 5 shows the Rule Type labels used by Fiesler et al. (2018), while Table 6 presents the new labels that we produced via an iterative, inductive coding process. The tables show the percentage of rules in the set that were tagged with each label.
Comparing our results to those of Fiesler et al. (Fiesler et al., 2018) illuminates some interesting differences between AI rules and general subreddit rules. Fiesler et al.’s results found that with general rules, Prescriptive rules, or rules that tell users what to do, occurred in about equal numbers as Restrictive rules that tell users what not to do [^7] These labels are not mutually exclusive: a rule could be tagged as both Prescriptive and Restrictive if it contains elements of both style.. However, in our AI Rules Set, $96.2\%$ of rules were tagged as Restrictive while only $79\%$ were tagged as Prescriptive. This further bears out in the data for our AI Rule: Stance label, which shows Unqualified Bans (e.g., “No AI-Generated Content”) as the most common label ($43.7\%$ of AI rules). Still, it is interesting to see that some communities are adopting stances other than outright bans, like the $32.2\%$ of AI rules labeled Stance: Qualified Ban and the $24.4\%$ labeled Stance: Disclosure. Perhaps this implies that communities will be flexible towards figuring out how to work with, rather than against, this new technology.
Overall, we observe that a high number of AI rules are related to images. Table 5 shows that $50\%$ of rules were tagged with the Images label from Fiesler et al. (2018), which according to our codebook, indicates that the rule “is about posting visual content like images, photos, or paintings”. For comparison, Fiesler et al. (2018) tagged less than $8\%$ of the rules in their dataset with this label. We see slightly smaller numbers with our inductively coded labels, which used a codebook that looked for more specific language: Table 6 indicates that $36\%$ of AI rules are labeled as Medium: Images and $28\%$ are labeled Medium: Art. This observation further supports the finding in Section 4.2 that subreddits with Art and Celebrity topics are more likely to have AI rules, as these are usually image-based subreddits.
Tables 5 and  6 also demonstrate that many AI rules on Reddit focus on concerns about content quality. Around $35\%$ of rules were labeled as relating to Low quality content; $17\%$ and $11\%$ respectively were labeled as Rationale: Low quality and Rationale: Low effort, meaning that they explicitly stated these concerns as their rationale for regulating AI use. These rationales are more prevalent than rationales based on accuracy, authenticity, and ethics. The focus on quality may suggest that explicit AI regulation is driven by practical concerns about content and reputation, not ideological objections over the use of AI more broadly. Notably, almost half of the AI rules were labeled Stance: Unqualified ban, indicating that many subreddits may find it easier to simply ban the use of AI outright, rather than trying to develop policies that can accommodate it under certain circumstances (e.g., requiring the disclosure of AI use).
Table 5. Prevalence of Rule Type labels in the AI Rules Set ($n=4,150$). We show the percentage of rules in the set that were tagged with each label. Rules may be tagged with multiple labels and each subreddit may have multiple rules included in this set.
Rule Type %
Content and behavior 99.78
Restrictive 96.19
Prescriptive 78.96
Images 50.48
Consequences, moderation, and enforcement 38.96
Low quality content 34.6
Post format 29.45
Copyright and piracy 17.23
Links and outside content 11.20
Spam 8.99
Off topic 7.3
NSFW 4.65
Reposting 4.34
Advertising and commercialization 3.08
Reddiquette and sitewide rules 1.01
Politics 0.94
Harassment 0.75
Doxxing and personal information 0.55
Spoilers 0.48
Hate speech 0.39
Trolling 0.39
Voting 0.1
Personal army 0.05
Personality 0.02
Table 6. Prevalence of AI Rule Type labels in the AI Rules Set ($n=4,150$). We show the percentage of rules in the set that were tagged with each label. Rules may be tagged with multiple labels and each subreddit may have multiple rules included in this set. Examples are drawn from the data but have been shortened for brevity.
AI Rule Type Percentage of AI Rules (%) Example
Rationale Low quality 16.94 Low quality content will be removed. This includes AI generated images and other AI media.
Low effort 11.01 AI generated images or Chatgpt text content will be removed as “low effort” content. Please do not post AI generated content on this subreddit.
Inauthentic 10.34 NO FAKES OR OVERLY SHOPPED PHOTOS - may result in a ban, serious offense. Images must be posted as published except for Quality AI-enhanced images. We ask that those titles include ”Enhanced.”
Original content 5.25 AI generated text does not qualify as an original source.
Off topic 3.4 Content generated by a computer algorithm is off topic and unwanted.
Spam 3.13 Any form of spam / blogspam is not allowed. This includes links to book reviews, AI generated content, bots, social media channels and any affiliate links.
Ethics 2.12 Given Miyazaki’s vehement opposition to AI and the ethical issues with AI generated art, posts that feature content generated by an AI model will be removed.
Policy 1.47 Deep fakes violate reddit TOS. You will be banned if you post one.
Misleading 1.42 AI altered videos are misleading and will be removed.
Inaccurate 0.96 AI is awesome, but when used to answer technical questions it often gives incorrect or sometimes dangerous recommendations.
Inhuman 0.89 We believe Hip hop is human made, and for that reason we’ve decided to ban AI generated content.
Stance Unqualified ban 43.73 AI-generated content of any kind is forbidden.
Qualified ban 32.24 AI Generated Fan Art is also not allowed, UNLESS it has been substantially edited by a human.
Disclosure 24.36 If posting AI-generated art, specify in the title that it is AI-created, include what service used.
Enabling 24.29 Recipes written by AI are ok but you must: 1) make the salsa, and 2) post a real photo of the completed salsa.
Requiring 0.92 Only post images generated by Ai, that’s it.
Medium Content 35.76 On r/NFT, content that is fully generated by AI and not modified in any way shape or form is prohibited and will result in a ban from the community.
Images 35.73 Please no DALL·E mini or other AI generated images.
Art 27.93 No AI art allowed, as per our content policy.
Fakes 21.88 No fake images of Leah Willamson are permitted, either photoshopped or AI generated
Unspecified 5.78 No AI - https://www.gofundme.com/f/protecting-artists-from-ai-technologies
Text 5.23 This includes both AI generated images/artwork and text generated by tools such as ChatGPT.
Deepfake 5.08 Deepfakes or other types of fakes are not allowed here, you will be banned for posting them.
Bots 3.37 Karma-farming with ChatGPT bots, Malware, phishing, spam or scam will result in your account being banned without warning.
Videos 2.84 No low effort AI content or Videos.
Framing Generate 51.49 AI-generated content of any kind is forbidden.
Assist 23.69 Traditional paintings and drawings only. No videos, screenshots, photography, audio, multi-panel comics, content generated or AI-assisted submissions.
Unspecified 8.07 The removal of low quality and, AI submissions, are subject to remove at mod discretion.
Since the Art and Celebrity Topics featured prominently in Section 4.2, we also separately examined patterns among Rule Types within each of these Topics. Among subreddits in the Art Topic, we find a higher share of AI rules with the Copyright and Piracy label ($34\%$). Among Celebrity subreddits, we see that the Low Quality Content and Format Rule Types labels are much more common ($74\%$ and $71\%$). In terms of the explicit rationales given for the AI rules, almost half of all AI rules in Celebrity subreddits were labeled Rationale: Low Quality and $29\%$ were labeled Rationle: Inauthentic. Additionally, $86\%$ of AI rules from Celebrity subreddits explicitly govern the use of “Fakes” (labeled with Medium: Fakes), which upon closer examination appears to be referring to heavily-edited photos.
5. Discussion
Our quantitative study offers two main research contributions: (1) a large-scale analysis of rules about AI use in Reddit communities and (2) a taxonomy for classifying these rules. Our study is an effort to support community self-moderation inspired by Seering’s second “guiding question” for the HCI research community: “What are the processes of context-sensitivity in online community moderation, and how might they be better supported?” (Seering, 2020). The sudden arrival of generative AI poses unique challenges for online communities (Lloyd et al., 2023) that HCI research is uniquely positioned to help address. Below we put our findings in dialogue with prior work and discuss design implications for social platforms.
Our findings show that subreddit communities, especially large ones (in terms of number of subscribers), are changing their rules to respond to generative AI. The new rules focus on AI-produced images more than any other medium, and often cite concerns about quality. The likelihood of having rules about AI varies with the type of community, with communities related to art or celebrities more likely to have these rules, and communities that are oriented around social support less likely.
The differences between our results and those of Fiesler et al. (Fiesler et al., 2018) emphasize the importance of governing AI usage to image-sharing communities. Our findings that Restrictive AI rules are more common than Prescriptive ones seems to suggest that, at least at this point in time, communities appear more interested in clarifying how AI cannot be used than how it can. Still, there is some evidence of flexibility in these stances, and it will be interesting to observe if they grow more or less flexible over time. We reiterate Fiesler et al.’s point that rules on Reddit are context-sensitive, so it is hard to generalize about the situations when communities opt for disclosure rather than an outright ban, but digging into these cases could be a topic of future HCI research.
Evaluating our results in the context of the qualitative findings of Lloyd et al. (2023) paints a nuanced picture of the prevalence of some of the concerns raised in their interviews. For example, over $50\%$ of rules in our AI Rules Set are about images, but the majority of concerns raised in Lloyd et al. pertain primarily to text. This is a significant difference as helping communities respond to AI-generated text as opposed to AI-generated images could be a very different task. Another interesting difference is that we find the Social Support Community Archetype (Prinster et al., 2024) to be negatively correlated with the presence of AI rules, which goes against one of the theories presented by Lloyd et al., that communities who seek to encourage authentic human connection may object the most to AIGC. We do find concerns about authenticity, but it tends to be more about the authenticity of content, not of interpersonal communication. Our findings that Copyright and Piracy AI rules are more common in Art subreddits echoes some of the quotes from the moderators’ of art-related communities interviewed in Lloyd et al. who object to AI because of how current AI models have been trained. The prevalence of AI rules in Celebrity subreddits also echoes the concerns about authenticity raised in Lloyd et al.’s interviews. Rules from Celebrity subreddits tended to be about deepfakes or heavily altered images, which recent studies of AIGC in the wild have identified as the most common form of misuse (Marchal et al., 2024). Lloyd et al. also emphasizes a concern that AIGC is low-quality content, which is confirmed by our analysis to be the most universal explicit objection across the platform. What a community considers to be “low quality” is extremely context-dependent, but this could be one area where UI features that clearly communicate norms and expectations to community members, such as rules that display at post creation time, could go a long way to positively shape user behavior (Matias, 2019).
Looking ahead, it will be interesting to see how these rules—and norms—continue to change over time. So far our data shows that AI rules have become more prevalent, but it is possible that these rules could disappear if AI use becomes widespread to the point that it is unremarkable, or where bans are impossible to enforce. Communities who care about gate-keeping (Keegan and Gergle, 2010) may seek out other ways for users to demonstrate their authenticity, similar to those that have developed in some anonymous online communities (Bernstein et al., 2011). The HCI community should keep an eye on how these rules change over time in order to continue to support these communities as their norms and practices evolve.
Our study is limited in that it is based on a sample of available metadata for subreddits that were publicly listed on Reddit’s web interface in the summer of 2023. We recollected metadata for this sample one year later, but did not add any new subreddits, which means that new communities created since our initial collection are excluded from our study. By collecting only publicly listed subreddits for our sample, our sample is also missing subreddits that are public but unlisted. This makes it likely that certain types of subreddit are underrepresented, such as NSFW subreddits. Our sample also excluded subreddits with a primary language other than English. Future research could explore whether the trends that we observed still hold across the types of communities excluded from our sample.
Our study also limits our analysis to text that is contained within a subreddit’s ”rules” metadata field. There are various other ways that online communities communicate rules and norms, such as through public discussions, wikis, direct messages, or moderation actions. Future work interested in studying communities’ attitudes towards AI use could consider these other sources of data, as well as qualitative approaches that would elicit direct input from community members and moderators.
Ethics Statement. Our study was not required to undergo IRB review as we worked only with publicly available data. Still, we recognize the importance of conducting ethical social media research and have taken steps to follow the best practices of the HCI community (Fiesler et al., 2024). We did not collect any information about individual users or conversations. Additionally, we did not include in our study any subreddits who were not listed in Reddit’s public index of communities. We removed from our sample any subreddits who were public during our first crawl but no longer accessible during our second. Our paper does not reference any of the specific communities in our sample, only aggregate statistics. We plan to make our datasets public after this publication completes the peer-review process, but will take additional steps to ensure no sensitive information is contained before doing so.
6. Conclusion
Reddit communities across the platform have responded to the arrival of generative AI by updating their public rules to state explicit positions on the use of this new technology. A specific community’s attitude towards the use of AIGC is nuanced and context-dependent, but our findings suggest a general trend of opposition from communities who value quality and authenticity. It remains to be seen how these rules will be enforced, but HCI researchers should stay attuned to the changing needs and norms of these online communities as they continue to adapt to this new technology.
Acknowledgements.
Appendix A LLM Prompts
A.1. Detecting Community Rules Governing AI Use
1You will be given the public community rules for a subreddit. You will determine if any of the subreddits rules explicitly govern the use of AI tools or the content that such tools produce. Please return a JSON object with your determination, either YES or NO, and your explanation.
2
3Examples:
4
51. Input: - AI generated(AI generated ): No AI (chatGPT etc.) generated messages/comments. No questions about chatGPT/AI generated code.
6Output: {"determination": "YES", "explanation":"This rule bans code generated with ChatGPT or other AI tools"}
7
82. Input: - No bots(No bots): No bots. All bots are banned unless the bot provides exceptional value to /r/ MealPrepSunday.
9Output: {"determination": "NO", "explanation":"This rule bans bots, not AI. Bots are not necessarily AI."}
10
113. Input: - Tributes, Split Edit, Deepfakes, roleplay, and DM Request are not allowed(Tributes, Split Edit, Deepfakes, roleplay, and DM Request are not allowed): The mods will take strict action against members who didnt follow this rule. The users who post Tributes will suffer a temporary ban.
12Output: {"determination": "YES", "explanation":"This rule states that Deepfakes, which are a form of AI-generated content, are not allowed."}
13
144. Input: Rule 2 - Future Focus(Rule 2 - Future Focus): Submissions must be *future focused*. All posts must have an initial comment, a **Submission Statement**, that suggests a line of future-focused discussion for the topic posted. We want this submission statement to elaborate on the topic being posted and suggest how it might be discussed in relation to the future.
15AI-focused posts are only allowed on the weekend.
16Output: {"determination": "NO", "explanation":"This rule governs the use of posts about AI, not those made using it."}
17
185. Input: - Restricted accounts(New, Low Karma, bot, and underage accounts are restricted): * New or low karma accounts - may not post, can participate in the comments
19* Negative karma accounts - may not post or comment
20* Under 18 users - may not post or seek diet advice
21* No bot / AI accounts
22Output: {"determination": "YES", "explanation":"This rule bans AI accounts."}
23
246. Input: - Rule 1: Must be relevant to language models(Rule 1: Must be relevant to language models): All language models are welcome here.
25
26Dall-E, GPT-3, GPT-Neo, GPT-J, ...
27Output: {"determination": "NO", "explanation":"This rule says that posts must be about AI but says nothing about their use."}
28
297. Input: - No fakes or overly shopped photos, images must be posted as published. No memes.(No fakes or overly shopped photos, images must be posted as published. No memes.): NO FAKES OR OVERLY SHOPPED PHOTOS - may result in a ban, serious offense. Images must be posted as published except for Quality AI-enhanced images. We ask that those titles include ""Enhanced."" Cartoonish/amateurish attempts will be removed solely at the mods discretion. Dont complain. Do better. ! No X-rays or ""bubbling."" No watermarks. No emojis. No weird borders on the images (i.e., white or black bars). No text on images.
30Output: {"determination": "YES", "explanation":"This rule specifies that all AI-enhanced images must be titled Enhanced’"}
31
328. Input: Properly Source Fan Content or Concept Art(This fan/concept art isnt properly sourced): We like to support the community and creators of the Marvel Cinematic Universe as much as possible and properly crediting them for their work is the least we can do. When submitting fan art or a piece of concept art, we ask that you properly link to the original source and mention the name of the artist in the title.
33Output: {"determination": "NO", "explanation":"This rule does not mention art produced by AI."}
A.2. Applying Subreddit Classification Labels
1You are to assist with a task to qualitatively code Reddit communities, which are known as subreddits.
2You will be given a JSON string containing metadata about a subreddit that will contains its name, topic, public description, description and rules.
3For each category below, follow the instructions to determine which labels apply to the subreddit.
4Return a JSON object where they keys are "Topical Question and Answer", "Learning and Perspective Broadening", "Social Support", "Content Generation", "Affiliation with an Entity", "Topic" and the values are the most appropriate label values.
5
6Category: Community Archetype
7Instructions: An archetype describes an abstract set of qualities that are likely to co-occur in individual instances of an object. A Community Archetype applies this concept to subreddits and classifies them into types of communities based on identifiable subreddit features and common user behaviors. The following labels each correspond to a Community Archetype. Consider each label and the characteristics associated with it and determine if it describes the given subreddit. For each label, choose the value Y if it best describes the subreddit and N if it does not.
8Labels:
9- Name: Topical Question and Answer
10 Characteristics: Subreddits that are explicitly set up with strict rules to create a Question and Answer format around specific topics. Most top-level posts are questions, while most comments are answers, or discussions and enrichments of others answers.
11- Name: Learning and Perspective Broadening
12 Characteristics: Subreddits that users visit primarily to learn or broaden their perspective. Top-level posts are often pointers to current events, publications, or relevant news, or relatable personal stories, experiences, and questions. Comments tend to elaborate upon the ideas raised, for example by celebrating or disagreeing, making jokes, providing contrasting or similar personal stories and experiences, or adding new ideas and references into the mix.
13- Name: Social Support
14 Characteristics: Subreddits where socially supportive behaviors are the main purpose and organizing principle for gathering. Top-level posts are often specific personal experiences or sensitive disclosures, questions about a health issue, announcements of milestones, reflections or venting, or sharing of resources, artwork, and encouraging thoughts and memes/jokes. Comments generally offer support, reflection, commiseration, resources, and validation.
15- Name: Content Generation
16 Characteristics: Subreddits that are devoted to content with a specific format, sense of humor, viewpoint, purpose, or type of creative expression. Top-level posts exemplify a subreddits specific content style and comments often include peoples opinions on the content, extra information about the content, or sometimes commiseration with the original poster.
17- Affiliation with an Entity
18 Characteristics: Subreddits that have explicit affiliations with particular entities, such as geographical places or organizations (cities, universities), popular media (book or fan series, TV shows), sports teams, etc. Posts often feature local or breaking news and events related to the entity. In the comments, users generally express their feelings about news or events, and answer questions or offer advice and opinions about the entity.
19
20Category: Topic
21Instructions: Pick one of the following labels that best describes the topic of the subreddit. If multiple labels apply, choose the one that is most descriptive.
22Labels:
23- Video Games
24- Image Sharing
25- Entertainment
26- Personal
27- Technology and Science
28- Celebrity
29- Hobby
30- Local
31- Sports
32- Support
33- Music
34- Art
35- Humor
36- Meta
37- Business and Organizations
38- Animals
39- Work
40- Writing
41- Learning
42- Politics
43- Games
44- News
45- Health
46- Food
47- Parody
48- Culture
49- Fashion
50- Drugs
A.3. Applying Rule Classification Labels
1You will be given one of a subreddits public community rules and must determine which labels apply.
2Follow the instructions for each label and determine if the label applies to the rule.
3Return a JSON object where the keys are the label names and the values are TRUE if the label applies and FALSE if not.
4
5Category: rule types
6
7Label name: Advertising and Commercialization
8Instructions: Is this rule about advertising and commercialization in the community?
9
10Label name: Consequences/Moderation/Enforcement
11Instructions: Does this rule specify what will happen if the rule is broken?
12
13Label name: Content/Behavior
14Instructions: Does this rule govern the types of behavior or content that are allowed or prohibited in the subreddit?
15
16Label name: Copyright/Piracy
17Instructions: Does this rule have to do with copyright, piracy, or giving credit to contents original creator?
18
19Label name: Doxxing/Personal Info
20Instructions: Is this rule about doxxing or revealing someones personal information?
21
22Label name: Format
23Instructions: Does this rule specify that posts must follow a specific format?
24
25Label name: Harassment
26Instructions: Is this rule about harassment?
27
28Label name: Hate Speech
29Instructions: Is this rule about hate speech?
30
31Label name: Images
32Instructions: Is this rule about posting visual content like images, photos, or paintings?
33
34Label name: Links and Outside Content
35Instructions: Is this rule about posting links or outside content?
36
37Label name: Low-Quality Content
38Instructions: Is this rule about posting low-quality content?
39
40Label name: NSFW
41Instructions: Is this rule about posting NSFW content?
42
43Label name: Off-topic
44Instructions: Is this rule about posting off-topic content?
45
46Label name: Personal Army
47Instructions: Is this rule about using a group of other members to argue on your behalf?
48
49Label name: Personality
50Instructions: Does this rule govern what personality traits belong in the community?
51
52Label name: Politics
53Instructions: Does this rule govern posts about politics?
54
55Label name: Prescriptive
56Instructions: Does this rule include any positive statements about what users must do, or what is allowed?
57
58Label name: Reddiquette/Sitewide
59Instructions: Does this rule reference sitewide policy or norms?
60
61Label name: Reposting
62Instructions: Is this rule about reposting content that has already been posted?
63
64Label name: Restrictive
65Instructions: Does this rule include any negative statements about what users must not, or are not allowed to, do?
66
67Label name: Spam
68Instructions: Is this rule about posting spam, such as karma farming, or posting a large volume of content?
69
70Label name: Spoilers
71Instructions: Is this rule about posting spoilers?
72
73Label name: Trolling
74Instructions: Is this rule about trolling?
75
76Label name: Voting
77Instructions: Is this rule about voting in the community?
78
79Category: Framing
80Label name: framing--generate
81Instructions: Does this rule use variations of the words "create" or "generate" to discuss content made with AI, such as AI-generated content?
82
83Label name: framing--assist
84Instructions: Does this rule reference using AI assistance, for example to enhance, alter, or edit images?
85
86Label name: framing--unspecified
87Instructions: Does this rule mention AI, but does not mention a specific type of use?
88
89Category: Rationale
90Label name: rationale--lowquality
91Instructions: Does this rule explicitly mention that content made with AI is low quality?
92
93Label name: rationale--loweffort
94Instructions: Does this rule mention that content made with AI is low effort?
95
96Label name: rationale--spam
97Instructions: Does this rule mention that content made with AI is spam?
98
99Label name: rationale--misleading
100Instructions: Does this rule mention that content made with AI is misleading?
101
102Label name: rationale--inaccurate
103Instructions: Does this rule mention that content made with AI is inaccurate?
104
105Label name: rationale--inauthentic
106Instructions: Does this rule mention that content made with AI is fake or inauthentic?
107
108Label name: rationale--inhuman
109Instructions: Does this rule mention that content made with AI is inhuman?
110
 [^1] Label name: rationale--offtopic
112Instructions: Does this rule mention that content made with AI is off topic?
113
114Label name: rationale--policy
115Instructions: Does this rule mention laws or site policies?
116
117Label name: rationale--ethics
118Instructions: Does this rule mention an ethical or moral objection to AI?
119
120Label name: rationale--original_content
121Instructions: Does this rule mention that content made with AI is not original content?
122
123Category: Medium
124Label name: medium--text
125Instructions: Does this rule explicitly mention text made with AI?
126
127Label name: medium--images
128Instructions: Does this rule explicitly mention images generated or enhanced by AI?
129
130Label name: medium--videos
131Instructions: Does this rule explicitly mention videos made with AI?
132
133Label name: medium--art
134Instructions: Does this rule explicitly mention art made with AI?
135
136Label name: medium--deepfake
137Instructions: Does this rule explicitly mention deepfakes?
138
139Label name: medium--fakes
140Instructions: Does this rule explicitly mention fakes?
141
142Label name: medium--bots
143Instructions: Does this rule explicitly mention AI bots?
144
145Label name: medium--content
146Instructions: Does this rule explicitly mention AI content, for example AI-generated content?
147
148Label name: medium--unspecified
149Instructions: Does this rule govern the use of AI without explicitly referencing a specific type of use?
150
151Category: Stance
152Label name: stance--unqualified_ban
153Instructions: Does this rule ban or disallow AI or some sort of content made using AI without providing exceptions?
154
155Label name: stance--qualified_ban
156Instructions: Does this rule ban some forms of AI use, but allow it in certain circumstances or if users label it?
157
158Label name: stance--disclosure
159Instructions: Does this rule require that users identify when the content that they post is made to some extent using AI, such as with a label, or flair, or certain text in the title?
160
161Label name: stance--enabling
162Instructions: Does this rule explictly say that some content made with AI is allowed under certain circumstances, for example if the fact that AI was used is disclosed?
163
164Label name: stance--requiring
165Instructions: Does this rule explictly require that posted content must be made with AI?